QUESTION 31

What is the molar concentration (also known as the molarity) of propionic acid in a 8.3 %(m/v) propionic acid solution. The formula for propionic acid is CH3CH2COOH.

QUESTION 32

Britney and Kristy are throwing a Halloween party. Britney and Kristy make a fruit punch by mixing 465. mL of orange juice with 2.23 liters of cranberry juice. Their teacher sneaks into the party and adds 228. mL of ethanol to the punch before anyone drinks it. What is the concentration of ethanol in the punch using the %(v/v) method? Remember that there are
1,000 mL in 1 L.

QUESTION 33

Tamara has 107 g of a NaCl solution that has concentration of 4.5 % NaCl (m/m). How many grams of NaCl are dissolved in that solution?
Answer #1

31)

consider 1 L of solution

volume = 1 L = 1000 mL

Now use:

(m/v)% = mass*100/volume

8.3 = mass*100/1000 mL

mass = 83 g

Molar mass of CH3CH2COOH = 3*MM(C) + 6*MM(H) + 2*MM(O)

= 3*12.01 + 6*1.008 + 2*16.0

= 74.078 g/mol

mass of CH3CH2COOH = 83 g

we have below equation to be used:

number of mol of CH3CH2COOH,

n = mass of CH3CH2COOH/molar mass of CH3CH2COOH

=(83.0 g)/(74.078 g/mol)

= 1.12 mol

volume , V = 1 L

we have below equation to be used:

Molarity,

M = number of mol / volume in L

= 1.12/1

= 1.12 M

Answer: 1.12 M

Only 1 question at a time please